The questions below refer to the following situation:
In an Analysis of Variance with 3 groups, each containing 15 respondents:

Calculate the between-group degrees of freedom.

a. 2

b. 3

c. 20
Answer #1

Given that, in an Analysis of Variance with 3 groups, each containing 15 respondents:

Number of groups = k = 3

Degrees of freedom between groups = k - 1 = 3 - 1 = 2

=> Between-group degrees of freedom = 2

Answer : a) 2